Five times the smallest of three consecutive odd integers is nine more than four times the second integer. Find the third integer.
Answer by JimboP1977(311)

let x be the smallest of the three consecutive integer
5x - 9 = 4(x+2)
5x-9 = 4x+8
5x-17 = 4x
-17 = -x
17 = x
so the third integer is x+4 = 21.